OpenStudy (anonymous):

This problem always confuses me! f(x) = 4x+5 g(x) = -x^2-6x+8 Find: g(x+h) - g(x)/h

OpenStudy (yttrium):

Let me just give you one example. let f(x) = x^2 + 3x -4 when i say find f(x+h) it is simply as f(x+h) = (x+h)^2 + 3(x+h) - 4 i hope you understand now

OpenStudy (campbell_st):

ok... so the question is about 1st principles... g(x + h) = -(x + h)^2 - 6(x+h) + 8 so you would need to distribute and simplify... can you do that..?

OpenStudy (anonymous):

how would the distribution of g(x-h) look like. if I can see that then i will be able to understand

OpenStudy (yttrium):

take a look on my example f(x) = x^2 + 3x - 4 f(x-h) = (x-h)^2 + 3 (x-h) - 4

OpenStudy (anonymous):

ok that makes sense thnx

OpenStudy (yttrium):

so, show your answer here now. let us check if you really understood :D

OpenStudy (anonymous):

haha ok this is what i got -(h+6+2x)

OpenStudy (anonymous):

so is it right or wrong

OpenStudy (yttrium):

show your solution too :D

OpenStudy (anonymous):

as in my work?

OpenStudy (yttrium):

yup

OpenStudy (anonymous):

alright (( -(x+h)^2 - 6(x+h) + 8 )) + (x^2 +6x + 8)/ h :thats how i set up the problem and then i distributed, combined like terms and then i divided it all by h

OpenStudy (anonymous):

by the way some of the plus signs look really faint

OpenStudy (anonymous):

do u see anything wrong?

OpenStudy (yttrium):

well, i find it right.

OpenStudy (anonymous):

ok thnx once again